OpenStudy (jazzyfa30):

Pre cal help Change the equation to polar coordinate. y^2=9x

OpenStudy (jazzyfa30):

@amistre64

OpenStudy (amistre64):

hmmm, need to define it as its distance from the origin eh

OpenStudy (jazzyfa30):

I dont know

OpenStudy (amistre64):

well, that is what a polar does, for some degree rotation, we are this far from the origin

OpenStudy (jazzyfa30):

oh

OpenStudy (amistre64):

some identities we can apply are: y = r sin(t) x = r cos(t) r^2 = x^2 + y^2

OpenStudy (anonymous):

does this help?

OpenStudy (jazzyfa30):

what is that?

OpenStudy (jazzyfa30):

is that polar form cus I dont know

OpenStudy (amistre64):

y^2 = 9x x^2 + y^2 = x^2 + 9x r^2 = r^2 cos^2(t) + 9r cos(t) r^2 = r^2 cos^2(t) + 9r cos(t) r^2 - r^2 cos^2(t) - 9r cos(t) = 0 (1- cos^2(t)) r^2 - (9cos(t)) r = 0 using the quadratic formula ... maybe\[r=\frac{-9cos(t)\pm\sqrt{81cos^2(t)}}{2(1-cos^2(t))}\] \[r=\frac{-18cos(t)}{2sin^2(t)}~or~0\]

OpenStudy (jazzyfa30):

is that the answer if so can you explain it step by step

OpenStudy (amistre64):

or r = -9 cot(t) csc(t)

OpenStudy (amistre64):

i did explain it step by step :/

OpenStudy (jazzyfa30):

nevermind

OpenStudy (jazzyfa30):

Thank you

OpenStudy (amistre64):

i prolly made some minor error that ill ahve to dbl chk ,,, but its all right there. sub in identities and hash it out with the algebra

OpenStudy (amistre64):

ugh, yeah, i did a -(-b) instead of a -b .... so ignore that - and it should be fine

OpenStudy (amistre64):

@rachelmcclellan thats might graph it and be some properties of the parabola but it does not seem to convert it a polar representation

OpenStudy (jazzyfa30):

Is the last answer you gave me the answer r = -9 cot(t) csc(t)

OpenStudy (amistre64):

go thru my initial material ... i was working out a thought to refresh my memory ... ask me where it does wrong for you

OpenStudy (jazzyfa30):

I dont know how to do it so i wouldnt be able to tell u where u went wrong

OpenStudy (amistre64):

all i did was use the identities to create a quadratic equation, and then solved for r. thats the best way i can approach it off the top of my head.

OpenStudy (amistre64):

if you are not adept at quadratics ... then this question might just be too far advanced for you to accomplish

OpenStudy (mathmale):

Change the equation to polar coordinate. y^2=9x: amistre64 was correct in stating that rectangular coordinates can be changed to polar coordinates through the following substitutions: x=r cos theta y=r sin theta. You have y^2=9x . So, from y=r sin theta, you need to obtain y^2. How? You have 9x on the right. Just substitute r sin theta for x. Your turn.

OpenStudy (jazzyfa30):

so